What is the holding period return of the car?


Holding period and annual (investment) returns. Bohenick Classic Automobiles restores and rebuilds old classic cars. The company purchased and restored a classic 1957 Thunderbird convertible 5 years ago for $8,300.00. Today at auction, the car sold for $74,700.00. What are the holding period return and the annual return on this investment? What is the holding period return of the car? [ % (Round to two decimal places )


What is the holding period return of the car?
Answer #1

HPR= P1-P0/ P0

= (74700-8300)/8300

=800%

Annual return = (1+ HPR)^(1/n) -1

= (1+8)^(1/5) -1

=55.18%
Answer #2

SOLUTION :


Holding period (5 years) returni in %

= (74700.00 - 8300.00) / 8300.00 *100

= 800 % (ANSWER).


Let annual return be r (in decimals)


So,


800 /100 = (1 + r)^5 - 1

=> r = (1 + 8.00)^(1/5) - 1

=> r = 0.5518 = 55.18% 


So, annual return = 55.18% (ANSWER).

answered by: Tulsiram Garg